OpenStudy (isbella):

1. can some explain to me on the proof how A'B'+B'C'=A'C' goes to this step =[r^2/(DA*DB)]AB + [r^2/(DB*DC)]BC + [r^2/(DA*DC)]AC then it goes to step AB*CD+BC*AC=AC*BD

OpenStudy (loser66):

A'B'+B'C'=A'C' ok? Now, by distance formula in inversion center D, we have A'B'= r^2 AB/ (DA*DB) B'C'= r^2 BC/(DB*DC) A'C'=r^2 AC/(DA*DC)

OpenStudy (loser66):

Put all in, we have \[\dfrac{r^2 AB}{DA*DB}+\dfrac{r^2BC}{DB*DC}=\dfrac{r^2AC}{DA*DC} \]

OpenStudy (loser66):

as it suggests, just multiple both sides by \(\dfrac{DA*DB*DC}{r^2}\) , you get what you want.

OpenStudy (loser66):

Proof of A'B'= r^2 AB/(DA*DB): We know that the inversion in circle center D radius r gives us DA*DA'= r^2 and we have \(\triangle DAB\cong \triangle DB'A'\) Hence \(\dfrac{DA}{DB'}=\dfrac{DB}{DA'}\)\(\rightarrow DB'=\dfrac{DA*DA'} {DB}=\dfrac{r^2}{DB}\) Same argument in similar triangles, we have \(\dfrac{AB}{B'A'}=\dfrac{DA}{DB'}\)\(\rightarrow A'B'=\dfrac{AB*DB'}{DA}\) replace \(DB'\) above You have \(A'B'=\dfrac{AB*r^2}{DA*DB}\) proof end